Two forces of 15N and 8N are inclined at 600 to each other. Use calculation method to find the resultant of the forces.

1
Expert's answer
2021-01-19T14:45:10-0500

The resultant force

"\bf F=F_1+F_2"

"F_x=F_{1x}+F_{2x}=15+8\cos 60^\circ=19\:{\rm N}\\
F_y=F_{1y}+F_{2y}=0+8\sin 60^\circ=6.9\:\rm N"

"F=\sqrt{F_x^2+F_y^2}=\sqrt{19^2+6.9^2}=20\:\rm N"